OpenStudy (anonymous):

Multiple choice: The concentration of a substance is 3.5M. If you diluted the substance in water using a 5 ste, 10-fold dilution process, what would the final concentration of the substance be? a) 3.5 mM b) 350nM c) 35nM d) 350uM e) 35uM

OpenStudy (anonymous):

5 step*

OpenStudy (mos1635):

C1*V1=C2*10V1=> C2=c1/10 5 steps Cinitial*Vin=Cfinal*Vf (but Vf=10^5 V1) Cfinal=Cinitial*10^-5..........